# Revisiting Cross-Diffusion for Overcrowding Dispersal in Interacting Species System

### v1

@@ 3.1 Preliminaries
in a bounded domain arises in various applications, especially in biology, to represent the evolution of species as they strive to minimize their internal energy $\mathcal{E}(\rho),$ which depends inherently on the density $\rho.$ Since the works [26] and [20], it has been established that the PDE (3.25) subject to homogeneous Neuman boundary condition can be derived by employing $\mathcal{E}(\rho)=\frac{1}{m-1}\int_{\Omega}\rho^{m}$ in the context of $W_{2}-$Wasserstein distance by setting $W^{\tau}(\rho,\tilde{\rho})=\frac{1}{2\tau}\mathcal{W}_{2}(\rho,\tilde{\rho})^{2}$. It is well known that it can be also derived by utilizing the steepest descent algorithm for the internal energy $\mathcal{E}(\rho)=\frac{1}{m+1}\int_{\Omega}\rho^{m+1}$ in the context of $H^{-1}-$norm. This is achieved by setting $W^{\tau}(\rho,\tilde{\rho})=\frac{1}{2\tau}\|\rho-\tilde{\rho}\|_{H^{-1}}^{2}$ (cf. [6, 19]). These concepts can be expanded to accommodate some monotone nonlinearity $\eta(r)$ instead of $r^{m},$ both for $W_{2}-$Wasserstein distance and also $H^{-1}-$norm cf. [22, 11, 12, 21]) frameworks. These frameworks can be also extended to nonlinear diffusion operators (see [1] for the $W_{c}-$Wasserstein distance framework, and the recent work [19] for the the case a generalized dual Sobolev approach).
@@
While similar at first glance, analyzing the dynamics of two interacting species using gradient flow becomes challenging. This paper focuses on a system with two species $(k=1,2)$, represented by densities $\rho_{1},$ $\rho_{2},$ where the state space is $X=X_{1}\times X_{2}.$ To capture the overcrowding dispersal dynamic, we consider an internal energy function dependent on the sum of both densities ; i.e. $\mathcal{E}(\rho)=\tilde{\mathcal{E}}(\rho_{1}+\rho_{2}).$ We employ a proximal optimization algorithm to construct a sequence of density pairs $(\rho_{1},\rho_{2})$ starting from initial densities $(\rho_{01},\rho_{02})$. The algorithm iteratively updates the densities using the following steepest descent formula:
(3.26) $\rho_{i}=\hbox{argmin}_{\rho\in X}\Big{\{}\tilde{\mathcal{E}}(\rho_{1}+\rho_{2})+W^{\tau}_{1}(\rho_{1}^{i-1},\rho_{1})+W^{\tau}_{2}(\rho_{2}^{i-1},\rho_{2})\Big{\}}).$
Here, $W^{\tau}_{k}$ measures the work required for each species $k$ to move from state $\rho_{k}^{i-1}$ to state $\rho_{k},$ in a small time step $\tau>0.$ For PME nonlinear diffusion process, previous works (see [10, 22, 13]) suggest that when $W^{\tau}_{k}$ is defined by the $\mathcal{W}_{2}$-Wasserstein distance and $\mathcal{E}(\rho)=\frac{1}{m-1}\int_{\Omega}(\rho_{1}+\rho_{2})^{m}$, this approach leads to a cross-diffusion system:
@@
(3.27) $\left\{\begin{array}[]{ll}{\rho_{1}}_{t}-\frac{m}{m-1}\>\nabla\cdot\left(\rho_{1}\>\nabla(\rho_{1}+\rho_{2})^{m}\right)=0\\ \\ {\rho_{2}}_{t}-\frac{m}{m-1}\>\nabla\cdot\left(\rho_{2}\>\nabla(\rho_{1}+\rho_{2})^{m}\right)=0\end{array}\right.\quad\hbox{ in }Q.$
@@
This system, subject to homogeneous Neumann boundary conditions and initial data, describes how both species avoid overcrowding regions through hyperbolic dynamics like in (1.2). However, in this paper we deviate from the previous approach and proposes using the minimum flow problem to measure the work required for state transitions. We consider:
@@
$-\nabla\cdot(\sigma_{k}\>\omega)=\rho_{k}^{i-1}-\rho_{k},\quad\text{ in }\Omega,$
with potential boundary conditions for $\sigma_{k}\>\omega\cdot\nu$ on $\partial\Omega\setminus\Gamma_{N_{k}}.$ The goal of $I_{k}(\rho_{k}^{i-1},\rho_{k})$ is to find the most efficient ”traffic scheme” for transferring mass from $\rho_{k}^{i-1}$ to $\rho_{k}.$ This scheme considers both the quadratic work of the flow $\omega$ between densities $\rho_{k}^{i-1}$ and $\rho_{k},$ and the charge $g_{k}$ for $\phi\cdot\nu$ at at specific boundary locations $\Gamma_{D}.$ Since the balance equation, (3.29) captures the essence of mass transfer, the transfer fee effectively models the diffusion process with boundary conditions, enabling analysis of overcrowding dynamics for two interacting species. Interestingly, unlike the single-species case, we’ll see that taking for instance $\mathcal{E}(\rho)=\frac{1}{m+1}\int_{\Omega}(\rho_{1}+\rho_{2})^{m+1}$ this approach leads to a fundamentally different cross-diffusion system generalizing in some sense (3.27) (see Remark 2). This new system is parabolic in nature :
(3.30) $\partial_{t}\rho_{k}-\nabla\cdot(\sigma_{k}\>\nabla\eta_{k})=0,\quad\hbox{ for}k=1,2,\quad\hbox{ in }Q:=(0,T)\times\Omega,$
where the potential $\rho_{k}$ and $\eta_{k}$ satisfies the spatial delimitation of region conditions (1.3) and the state equation
$\eta_{1}\vee\eta_{2}=(\rho_{1}+\rho_{2})^{m}.$
Moreover, the approach incorporates non-homogeneous boundary conditions like those in equation (1.6). In the following, section we precise the assumptions and give the rigorous proofs for the results in a general case. Then we show how one can use this approach to prove the results of Section 2. We show also how to connect the approach with gradient flow in dual Sobolev space in the homogeneous case.
@@ 3.2 Notations
    $L^{2}_{+}(\Omega)$ denotes the set on nonegative Lebesgue integrable functions of $L^{2}(\Omega).$
@@
    $[H^{-1}_{\Gamma_{D}}(\Omega)]^{2}=H^{-1}_{\Gamma_{D_{1}}}(\Omega)\times H^{-1}_{\Gamma_{D_{2}}}(\Omega)$
@@
    $[H^{-1/2}(\Gamma_{D})]^{2}:=H^{-1/2}(\Gamma_{D_{1}})\times H^{-1/2}(\Gamma_{D_{2}}).$ Moreover, for any $h=(h_{1},h_{2})\in[H^{-1/2}(\Gamma_{D_{1}})]^{2}\times H^{-1/2}(\Gamma_{D_{2}})$ and $g=(g_{1},g_{2})\in[H^{1/2}_{00}(\Gamma_{D})]^{2}$ we use the notation
@@
$\sigma\>\phi=(\sigma_{1}\phi_{1},\sigma_{2}\>\phi_{2})$
@@
$\beta_{\alpha}[\rho]:=\beta(.,\alpha_{1}\>\rho_{1}+\alpha_{2}\>\rho_{2}),\quad\hbox{ for any }\rho\in[L^{2}(\Omega)]^{2}.$
@@ 3.3 Proximal minimization for cross-diffusion modeling
be given. We denote by $\tilde{g}$ the $[H^{1}(\Omega)]^{2}$ function such that $\tilde{g}=0$ on $\Gamma_{N}$ and $\tilde{g}=g$ on $\Gamma_{D}.$
To define the transition work rigorously, for any $\mu\in[L^{2}(\Omega)]^{2}$ and $\chi\in[L^{2}(\Omega)^{N}]^{2}$, we consider
@@
In some sense, for each $k=1,2,$ $\mathcal{A}_{\pi}^{\chi}[\mu]_{k}$ is the set of all mass fluxes which balance the mass $\mu_{k},$ under the action of the external force $\chi_{k}$ and the boundary action $\pi_{k}$ on $\Gamma_{k}.$ Recall that (see for instance [19]) that, for any $k=1,2,$ we have
@@
$\mathcal{N}_{\pi,g}^{\chi}[\mu]:=\inf_{\rho,\phi}\left\{\int_{\Omega}\beta_{\alpha}[\rho]+\int_{\Omega}F_{\sigma}[\phi]-[(\sigma\>\phi+\chi)\cdot\nu,g]_{\Gamma_{D}}\>:\>\rho\in[L^{2}_{+}(\Omega)]^{2},\>\phi\in\mathcal{A}_{\pi}^{\chi}[\mu-\rho]\right\},$
@@
$\mathcal{N}_{\pi,g}^{\chi}[\mu]=\inf_{\rho,\omega}\left\{\int_{\Omega}\beta_{\alpha}[\rho]+\underbrace{\mathcal{I}_{\pi,g}^{\chi}[\mu-\rho]_{1}+\mathcal{I}_{\pi,g}^{\chi}[\mu-\rho]_{2}}_{\mathcal{I}_{\pi,g}^{\chi}[\mu-\rho]}\>:\>\rho\in[L^{2}_{+}(\Omega)]^{2}\right\},$
@@
(3.33) $\mathcal{I}_{\pi,g}^{\chi}[\mu]_{k}:=\inf_{\omega}\left\{\frac{1}{2}\int_{\Omega}\sigma_{k}\>|\omega|^{2}-\langle(\sigma\>\omega+\chi_{k})\cdot\nu,g_{k}\rangle_{\Gamma_{D_{k}}}\>:\>\omega\in\mathcal{A}_{\pi}^{\chi}[\mu]_{k}\right\}.$
Here the quantity $\mathcal{I}_{\pi,g}^{\chi}(\mu-\rho)$ represents the transition work associated with the distribution of mass $\mu-\rho$, and the term $\nabla\cdot\chi$ can be utilized to represent a significant external influence exerted on the system by an external force, $\chi$ (along with boundary conditions).
@@
(3.34) $\mathcal{I}_{\pi,g}^{\chi}[\mu]_{k}=\max_{z\in H^{1}(\Omega),\>z_{/\Gamma_{D_{k}}}=g_{k}}\left\{\int_{\Omega}\mu_{k}\>z-\int_{\Omega}\chi_{k}\cdot\nabla z-\ \int_{\Omega}F_{\sigma}[\nabla z]+\langle\pi_{k},z\rangle_{\Gamma_{N_{k}}}\right\}.$
@@
(3.36) $\mathcal{N}_{\pi,g}^{\chi}[\mu]=\underbrace{\max_{\eta\in[H^{1}(\Omega)]^{2},\eta_{/\Gamma_{D}}=g}\left\{\int_{\Omega}\mu\cdot\eta+\int_{\Omega}[\chi\cdot\nabla\eta]-\int_{\Omega}F_{\sigma}[\nabla\eta]-\int_{\Omega}\beta_{\alpha}^{*}[\eta]+[\pi,\eta]_{\Gamma_{N}}\right\}}_{=:D_{\pi,g}^{\chi}[\mu]},$
@@
$\beta_{\alpha}^{*}[\eta]:=\beta^{*}\left(x,\frac{\eta_{1}}{\alpha_{1}}\vee\frac{\eta_{2}}{\alpha_{2}}\right),\quad\hbox{ for any }\eta\in[L^{2}(\Omega)]^{2}.$
Moreover, $(\rho,\phi)$ and $\eta$ are solutions of $\mathcal{N}_{\pi,g}^{\chi}[\mu]$ and $D_{\pi,g}^{\chi}[\mu],$ respectively, if and only if for each $k=1,2,$ $\phi_{k}=\nabla\eta_{k}$ and the couple $(\rho,\eta)$ satisfies the following system of PDE :
(3.37) $\left\{\begin{array}[]{ll}\left.\begin{array}[]{ll}\rho_{k}-\nabla\cdot(\sigma_{k}\>\nabla\eta_{k})=\mu_{k}+\nabla\cdot\chi_{k},\quad\hbox{ for }k=1,2\\ \\ \rho_{1},\>\rho_{2}\geq 0,\>\frac{\eta_{1}}{\alpha_{1}}\vee\frac{\eta_{2}}{\alpha_{2}}\in\partial\beta(x,\alpha_{1}\rho_{1}+\alpha_{2}\rho_{2})\\ \\ \rho_{1}\left(\frac{\eta_{2}}{\alpha_{2}}-\frac{\eta_{1}}{\alpha_{1}}\right)^{+}=\rho_{2}\left(\frac{\eta_{1}}{\alpha_{1}}-\frac{\eta_{2}}{\alpha_{2}}\right)^{+}=0\\ \end{array}\right\}&\hbox{ in }\Omega,\\ \\ (\sigma_{k}\>\nabla\eta_{k}+\chi_{k})\cdot\nu=\pi_{k}&\hbox{ on }\Gamma_{N_{k}}\\ \\ \eta_{k}=g_{k}&\hbox{ on }\Gamma_{D_{k}},\end{array}\right.$
in the sense that $\rho_{k}\in L^{2}_{+}(\Omega),$ $\eta_{k}\in H^{1}(\Omega),$ $\eta_{k}=g_{k}$ on $\Gamma_{D_{k}},$ for $k=1,2,$ and we have
  • $\rho_{1}\left(\frac{\eta_{2}}{\alpha_{2}}-\frac{\eta_{1}}{\alpha_{1}}\right)^{+}=\rho_{2}\left(\frac{\eta_{1}}{\alpha_{1}}-\frac{\eta_{2}}{\alpha_{2}}\right)^{+}=0,$ a.e. in $\Omega,$
  • $\frac{\eta_{1}}{\alpha_{1}}\vee\frac{\eta_{2}}{\alpha_{2}}\in\partial\beta(x,\alpha_{1}\rho_{1}+\alpha_{2}\rho_{2})$, a.e. in $\Omega,$
  • For each $k=1,2,$
    $\int_{\Omega}\rho_{k}\>\xi_{k}+\int_{\Omega}(\sigma_{k}\>\nabla\eta_{k}+\chi_{k})\cdot\nabla\xi_{k}=\int_{\Omega}\mu_{k}\>\xi_{k}+\langle\pi_{k},\xi_{k}\rangle_{\Gamma_{k}}.$
    for any $\xi_{k}\in H^{1}_{\Gamma_{{}_{k}}}(\Omega).$
@@
$K[f]:=\int_{\Omega}f_{0}\cdot\tilde{g}-\int_{\Omega}[\overline{f}\cdot\nabla\tilde{g}]+\inf_{\rho,\phi}\left\{\int_{\Omega}\beta_{\alpha}[\rho]+\int_{\Omega}F_{\sigma}[\phi]\>dx-[(\sigma\>\phi+\chi+\overline{f})\cdot\nu,g]_{\Gamma_{D}}\>\right.$
$\left.:\>\rho\in[L^{2}_{+}(\Omega)]^{2},\>\phi\in\mathcal{A}_{\pi}^{\chi+\overline{f}}(\mu_{k}+f_{0}-\rho)\right\},$
where $f_{0}=(f_{01},f_{02})$ and $\overline{f}=(\overline{f}_{1},\overline{f}_{2})$ are given by the decomposition $f_{k}=f_{0k}+\nabla\cdot\overline{f}_{k}$ in $H^{-1}_{\Gamma_{D_{k}}}(\Omega),$ for each $k=1,2.$ Thank to (2.12), we see that
@@
(3.38) $\begin{array}[]{c}K[f]=\inf_{\rho,\phi}\left\{\int_{\Omega}\beta_{\alpha}[\rho]+\int_{\Omega}F_{\sigma}[\phi]\>dx-\int_{\Omega}[(\sigma\>\phi+\chi)\cdot\nabla\tilde{g}]+\int_{\Omega}(\mu-\rho)\cdot\tilde{g}\>\right.\\ \\ \>\left.:\>\rho\in[L^{2}_{+}(\Omega)]^{2},\>\phi\in\mathcal{A}_{\pi}^{\chi+\overline{f}}[\mu+f_{0}-\rho]\right\}.\end{array}$
The following duality result is the key to proving Theorem 3. After, we need simply to compute the Legendre transform of $K$ to complete the proofs.
@@
$\int\beta_{\alpha}[\rho_{t}]+\>\int_{\Omega}F_{\sigma}[\phi_{t}]-[(\sigma\>\phi_{t}+\chi+\overline{f}_{t})\cdot\nu,g]_{\Gamma_{D}}\leq t\left(\int_{\Omega}\beta_{\alpha}[\rho_{h}]+\>\int_{\Omega}F_{\sigma}[\phi_{h}]-[\phi_{h}+\chi+\overline{h})\cdot\nu,g]_{\Gamma_{D}}\right)$
$+(1-t)\left(\int_{\Omega}\beta_{\alpha}[\rho_{f}]+\>\int_{\Omega}F_{\sigma}[\phi_{f}]-[(\sigma\>\phi_{f}+\chi+\overline{f})\cdot\nu,g]_{\Gamma_{D}}\right),$
@@
$\displaystyle K[f]$$\leq$ $\displaystyle\int_{\Omega}\beta_{\alpha}[\rho]+\int_{\Omega}F_{\sigma}[\phi]-\int_{\Omega}[(\sigma\>\phi+\chi)\cdot\nabla\tilde{g}]+\int_{\Omega}(\mu-\rho)\cdot\tilde{g}$
@@
Thus the result. $\square$

### v2

@@ 3.1 Preliminaries
in a bounded domain arises in various applications, especially in biology, to represent the evolution of species as they strive to minimize their internal energy $\mathcal{E}(\rho),$ which depends inherently on the density $\rho.$ Since the works [26] and [19], it has been established that the PDE (3.36) subject to homogeneous Neuman boundary condition can be derived by employing $\mathcal{E}(\rho)=\frac{1}{m-1}\int_{\Omega}\rho^{m}$ in the context of $W_{2}-$Wasserstein distance by setting $W^{\tau}(\rho,\tilde{\rho})=\frac{1}{2\tau}\mathcal{W}_{2}(\rho,\tilde{\rho})^{2}$. It is well known that it can be also derived by utilizing the steepest descent algorithm for the internal energy $\mathcal{E}(\rho)=\frac{1}{m+1}\int_{\Omega}\rho^{m+1}$ in the context of $H^{-1}-$norm. This is achieved by setting $W^{\tau}(\rho,\tilde{\rho})=\frac{1}{2\tau}\|\rho-\tilde{\rho}\|_{H^{-1}}^{2}$ (cf. [6, 18]). These concepts can be expanded to accommodate some monotone nonlinearity $\eta(r)$ instead of $r^{m},$ both for $W_{2}-$Wasserstein distance and also $H^{-1}-$norm frameworks (cf. [21, 12, 13, 20]). These frameworks can be also extended to nonlinear diffusion operators (see [1] for the $W_{c}-$Wasserstein distance framework, and the recent work [18] for the the case a generalized dual Sobolev approach).
@@
This paper focuses on a system with two species $(k=1,2)$, represented by a couple of densities, $(\rho_{1},\rho_{2}),$ in the state space $X=X_{1}\times X_{2}.$ To capture the overcrowding dispersal dynamic, we consider an internal energy function dependent on the sum of both densities ; i.e. $\mathcal{E}(\rho)=\tilde{\mathcal{E}}(\rho_{1}+\rho_{2}).$ We employ a proximal optimization algorithm to construct a sequence of density pairs $(\rho_{1},\rho_{2})$ starting from initial densities $(\rho_{01},\rho_{02})$. The algorithm iteratively updates the densities using the following steepest descent formula:
(3.37) $\rho^{i}=\hbox{argmin}_{\rho\in X}\Big{\{}\tilde{\mathcal{E}}(\rho_{1}+\rho_{2})+W^{\tau}_{1}(\rho_{1}^{i-1},\rho_{1})+W^{\tau}_{2}(\rho_{2}^{i-1},\rho_{2})\Big{\}}.$
Here, $W^{\tau}_{k}$ measures the work required for each species $k$ to move from state $\rho_{k}^{i-1}$ to state $\rho_{k},$ in a small time step $\tau>0.$ For PME nonlinear diffusion process, the work in [21] suggests to take $W^{\tau}_{k}$ defined by the $\mathcal{W}_{2}$-Wasserstein distance and $\mathcal{E}(\rho)=\frac{1}{m-1}\int_{\Omega}(\rho_{1}+\rho_{2})^{m},$ where we assume to simplify the presentation that the drift is equal to $0$ This approach leads to interesting optimal transportation interpretation of the solution at the discrete time level using Kantorovich potentials, and transports maps for each specie. However, the continuum limit seems to be complex, and there does not seem to be a PDE approach to yield well-posedness on the continuum PDE, even if the authors seem to expect to cover with this approach (at least for the particular situations where the solutions are well-behaved to stay separated throughout the evolution, with stable interface in between them) the solutions of the cross-diffusion system :
@@
(3.38) $\left\{\begin{array}[]{ll}\partial_{t}{\rho_{1}}-\frac{m}{m-1}\>\nabla\cdot\left(\rho_{1}\>\nabla(\rho_{1}+\rho_{2})^{m}\right)=0\\ \\ \partial_{t}{\rho_{2}}-\frac{m}{m-1}\>\nabla\cdot\left(\rho_{2}\>\nabla(\rho_{1}+\rho_{2})^{m}\right)=0\end{array}\right.\quad\hbox{ in }(0,\infty)\times\Omega.$
@@
This system, subject to homogeneous Neumann boundary conditions and initial data, describes how both species avoid overcrowding regions through hyperbolic dynamics in the spirit of (1.2). However, in this paper we deviate from the previous approach and proposes using the minimum flow problem to measure the work required for state transitions. We consider:
@@
(3.41) $-\nabla\cdot(\sigma_{k}\>\omega)=\rho_{k}^{i-1}-\rho_{k},\quad\text{ in }\Omega,$
with possible boundary conditions for $\sigma_{k}\>\omega\cdot\nu$ on $\partial\Omega\setminus\Gamma_{N_{k}}.$ The goal of $I_{k}(\rho_{k}^{i-1},\rho_{k})$ is to find the most efficient ”traffic scheme” for transferring mass from $\rho_{k}^{i-1}$ to $\rho_{k}.$ This scheme considers both the quadratic work of the flow $\omega$ between densities $\rho_{k}^{i-1}$ and $\rho_{k},$ and the charge $g_{k}$ for $\phi\cdot\nu$ at at specific boundary locations $\Gamma_{D_{k}}.$ Since the balance equation, (3.40) captures the essence of mass transfer, the transfer fee effectively models the diffusion process with boundary conditions, enabling analysis of overcrowding dynamics for two interacting species. Interestingly, unlike the single-species case, we’ll see that taking for instance $\mathcal{E}(\rho)=\frac{1}{m+1}\int_{\Omega}(\rho_{1}+\rho_{2})^{m+1}$ this approach leads to a fundamentally different cross-diffusion system generalizing in some sense (3.38) (see Remark 2). This new system is parabolic in nature :
(3.42) $\partial_{t}\rho_{k}-\nabla\cdot(\sigma_{k}\>\nabla\eta_{k})=0,\quad\hbox{ for}k=1,2,\quad\hbox{ in }(0,\infty)\times\Omega,$
where the potential $\eta_{k}$ satisfies the state equations
(3.43) $\eta_{k}-(\rho_{1}+\rho_{2})^{m}\in\partial I\!\!I_{[0,\infty)}(\rho_{k}),\quad\hbox{ for each }k=1,2.$
Thanks to Remark 4, this is clearly connected to the maximization problem
$\max_{\rho_{1},\rho_{2}\geq 0}\Big{\{}\rho_{1}\eta_{1}+\rho_{2}\eta_{2}-\frac{1}{m+1}(\rho_{1}+\rho_{2})^{m+1}\Big{\}}.$
Moreover, the approach incorporates mixed boundary condition for each specie, respect to the choice of the boundary condition in (3.40) and (3.41).
In the following section, we precise the assumptions and give the rigorous proofs for the results in a general case. Then, we show how one can use this approach to prove the results of Section 2. We show also how to connect the approach with gradient flow in dual Sobolev space in the homogeneous case.
@@ 3.2 Notations
    $[H^{-1}_{\Gamma_{D}}(\Omega)]^{2}=H^{-1}_{\Gamma_{D_{1}}}(\Omega)\times H^{-1}_{\Gamma_{D_{2}}}(\Omega).$
@@
    $[H^{-1/2}(\Gamma_{D})]^{2}:=H^{-1/2}(\Gamma_{D_{1}})\times H^{-1/2}(\Gamma_{D_{2}}).$ Moreover, for any $h=(h_{1},h_{2})\in[H^{-1/2}(\Gamma_{D})]^{2}$ and $g=(g_{1},g_{2})\in[H^{1/2}_{00}(\Gamma_{D})]^{2},$ we use the notation
@@
$\sigma\>\phi:=(\sigma_{1}\phi_{1},\sigma_{2}\>\phi_{2})$
@@
$\beta[\rho]:=\beta(.,\rho_{1}+\rho_{2}),\quad\hbox{ for any }\rho\in[L^{2}(\Omega)]^{2}.$
@@ 3.3 Proximal minimization for cross-diffusion modeling
be given. We denote by $\tilde{g}$ the $[H^{1}(\Omega)]^{2}$ function such that
(3.46) $\tilde{g}=0\hbox{ on }\Gamma_{N}\hbox{ and }\tilde{g}=g\hbox{ on }\Gamma_{D}.$
To define rigorously the transition work, for any $\mu\in[L^{2}(\Omega)]^{2}$ and $\chi\in[L^{2}(\Omega)^{N}]^{2}$, we consider
@@
In some sense, for each $k=1,2,$ $\mathcal{A}_{\pi}^{\chi}[\mu]_{k}$ is the set of all mass fluxes which balance the mass $\mu_{k},$ under the action of the external force $\chi_{k}$ and the boundary action $\pi_{k}$ on $\Gamma_{k}.$ Recall that (see for instance [18]), for any $k=1,2,$ we have
@@
$\mathcal{N}_{\pi,g}^{\chi}[\mu]:=\inf_{\rho,\phi}\left\{\int_{\Omega}\beta[\rho]+\int_{\Omega}F_{\sigma}[\phi]-[(\sigma\>\phi+\chi)\cdot\nu,g]_{\Gamma_{D}}\>:\>\rho\in[L^{2}_{+}(\Omega)]^{2},\>\phi\in\mathcal{A}_{\pi}^{\chi}[\mu-\rho]\right\},$
@@
$\mathcal{N}_{\pi,g}^{\chi}[\mu]=\inf_{\rho,\omega}\left\{\int_{\Omega}\beta[\rho]+\underbrace{\mathcal{I}_{\pi,g}^{\chi}[\mu-\rho]_{1}+\mathcal{I}_{\pi,g}^{\chi}[\mu-\rho]_{2}}_{\mathcal{I}_{\pi,g}^{\chi}[\mu-\rho]}\>:\>\rho\in[L^{2}_{+}(\Omega)]^{2}\right\},$
@@
(3.47) $\mathcal{I}_{\pi,g}^{\chi}[\mu-\rho]_{k}:=\inf_{\omega}\left\{\frac{1}{2}\int_{\Omega}\sigma_{k}\>|\omega|^{2}-\langle(\sigma\>\omega+\chi_{k})\cdot\nu,g_{k}\rangle_{\Gamma_{D_{k}}}\>:\>\omega\in\mathcal{A}_{\pi}^{\chi}[\mu-\rho]_{k}\right\}.$
Here the quantity $\mathcal{I}_{\pi,g}^{\chi}[\mu-\rho]$ represents the transition work associated with the distribution of mass $\mu-\rho$, and the term $\nabla\cdot\chi$ can be utilized to represent a significant external influence exerted on the system by an external force, $\chi$ (along with boundary conditions).
@@
(3.48) $\mathcal{I}_{\pi,g}^{\chi}[\mu]_{k}=\max_{z\in H^{1}(\Omega),\>z_{/\Gamma_{D_{k}}}=g_{k}}\left\{\int_{\Omega}\mu_{k}\>z-\int_{\Omega}\chi_{k}\cdot\nabla z-\frac{1}{2}\int_{\Omega}\sigma_{k}\>|\nabla z|^{2}+\langle\pi_{k},z\rangle_{\Gamma_{N_{k}}}\right\}.$
@@
(3.50) $\mathcal{N}_{\pi,g}^{\chi}[\mu]=\underbrace{\max_{\eta\in[H^{1}(\Omega)]^{2},\eta_{/\Gamma_{D}}=g}\left\{\int_{\Omega}\mu\cdot\eta+\int_{\Omega}[\chi\cdot\nabla\eta]-\int_{\Omega}F_{\sigma}[\nabla\eta]-\int_{\Omega}\beta^{*}[\eta]+[\pi,\eta]_{\Gamma_{N}}\right\}}_{=:D_{\pi,g}^{\chi}[\mu]},$
@@
$\beta^{*}[\eta]:=\beta^{*}\left(x,\eta_{1}\vee\eta_{2}\right),\quad\hbox{ for any }\eta\in[L^{2}(\Omega)]^{2}.$
Moreover, $(\rho,\phi)$ and $\eta$ are solutions of $\mathcal{N}_{\pi,g}^{\chi}[\mu]$ and $D_{\pi,g}^{\chi}[\mu],$ respectively, if and only if $\phi_{k}=\nabla\eta_{k}$ and the couple $(\rho_{k},\eta_{k})$ satisfies the following system of PDE :
(3.51) $\left\{\begin{array}[]{ll}\left.\begin{array}[]{ll}\rho_{k}-\nabla\cdot(\sigma_{k}\>\nabla\eta_{k})=\mu_{k}+\nabla\cdot\chi_{k}\\ \\ \eta_{k}\in\partial\beta(.,\rho_{1}+\rho_{2})+\partial I\!\!I_{[0,\infty)}(\rho_{k})\end{array}\right\}&\hbox{ in }\Omega,\\ \\ (\sigma_{k}\>\nabla\eta_{k}+\chi_{k})\cdot\nu=\pi_{k}&\hbox{ on }\Gamma_{N_{k}}\\ \\ \eta_{k}=g_{k}&\hbox{ on }\Gamma_{D_{k}},\end{array}\right.\quad\hbox{ for }k=1,2$
in the sense that $\rho_{k}\in L^{2}_{+}(\Omega),$ $\eta_{k}\in H^{1}(\Omega),$ $\eta_{k}=g_{k}$ on $\Gamma_{D_{k}},$
$\eta_{1}\vee\eta_{2}=:\tilde{\eta}\in\partial\beta(x,\rho_{1}+\rho_{2}),\>\quad\eta_{k}-\tilde{\eta}\in\partial I\!\!I_{[0,\infty)}(\rho_{k}),\quad\hbox{ a.e. in }\Omega,$
and
$\int_{\Omega}\rho_{k}\>\xi_{k}+\int_{\Omega}(\sigma_{k}\>\nabla\eta_{k}+\chi_{k})\cdot\nabla\xi_{k}=\int_{\Omega}\mu_{k}\>\xi_{k}+\langle\pi_{k},\xi_{k}\rangle_{\Gamma_{k}},\quad\forall\>\xi_{k}\in H^{1}_{\Gamma_{{}_{k}}}(\Omega).$
@@
$K[f]:=\int_{\Omega}f_{0}\cdot\tilde{g}-\int_{\Omega}[\overline{f}\cdot\nabla\tilde{g}]+\inf_{\rho,\phi}\left\{\int_{\Omega}\beta[\rho]+\int_{\Omega}F_{\sigma}[\phi]\>dx-[(\sigma\>\phi+\chi+\overline{f})\cdot\nu,g]_{\Gamma_{D}}\>\right.$
$\left.:\>\rho\in[L^{2}_{+}(\Omega)]^{2},\>\phi\in\mathcal{A}_{\pi}^{\chi+\overline{f}}[\mu+f_{0}-\rho]\right\},$
where $f_{0}=(f_{01},f_{02})$ and $\overline{f}=(\overline{f}_{1},\overline{f}_{2})$ are given by the decomposition $f_{k}=f_{0k}+\nabla\cdot\overline{f}_{k}$ in $H^{-1}_{\Gamma_{D_{k}}}(\Omega),$ for each $k=1,2.$ Then, since $\mathcal{N}_{\pi,g}^{\chi}[\mu]=K[0],$ it is enough to prove that $K[0]$ coincides with $D_{\pi,g}^{\chi}[\mu].$ To this aim we use duality techniques which involves the Legendre transforms of $K.$ This is the aim of the following lemma.
Remark 5
Thank to (2.14), we see that $K$ may be rewritten by using $\tilde{g}$ given by (3.46). Indeed, thanks to (3.46), we have
@@
(3.52) $\begin{array}[]{c}K[f]=\inf_{\rho,\phi}\left\{\int_{\Omega}\beta[\rho]+\int_{\Omega}F_{\sigma}[\phi]\>dx-\int_{\Omega}[(\sigma\>\phi+\chi)\cdot\nabla\tilde{g}]+\int_{\Omega}(\mu-\rho)\cdot\tilde{g}\>\right.\\ \\ \>\left.:\>\rho\in[L^{2}_{+}(\Omega)]^{2},\>\phi\in\mathcal{A}_{\pi}^{\chi+\overline{f}}[\mu+f_{0}-\rho]\right\}.\end{array}$
@@
$\int\beta[\rho_{t}]+\>\int_{\Omega}F_{\sigma}[\phi_{t}]-[(\sigma\>\phi_{t}+\chi+\overline{f}_{t})\cdot\nu,g]_{\Gamma_{D}}\leq t\left(\int_{\Omega}\beta[\rho_{h}]+\>\int_{\Omega}F_{\sigma}[\phi_{h}]-[\phi_{h}+\chi+\overline{h})\cdot\nu,g]_{\Gamma_{D}}\right)$
$+(1-t)\left(\int_{\Omega}\beta[\rho_{f}]+\>\int_{\Omega}F_{\sigma}[\phi_{f}]-[(\sigma\>\phi_{f}+\chi+\overline{f})\cdot\nu,g]_{\Gamma_{D}}\right),$
@@
$\displaystyle K[f]$$\leq$ $\displaystyle\int_{\Omega}\beta[\rho]+\int_{\Omega}F_{\sigma}[\phi]-\int_{\Omega}[(\sigma\>\phi+\chi)\cdot\nabla\tilde{g}]+\int_{\Omega}(\mu-\rho)\cdot\tilde{g}$
@@
Thus the result.
$\square$